# Free multiderivations of connected subgraph arrangements

### v2

@@ Abstract.
Cuntz and Kühne [CK24] introduced the class of connected subgraph arrangements ${\mathscr{A}}_{G}$, depending on a graph $G$, and classified all graphs $G$ such that the corresponding arrangement ${\mathscr{A}}_{G}$ is free. We extend their result to the multiarrangement case and classify all graphs $G$ for which the corresponding arrangement ${\mathscr{A}}_{G}$ supports some multiplicity $\mu$ such that the multiarrangement $({\mathscr{A}}_{G},\mu)$ is free.
@@ 1. Introduction
The study of modules of logarithmic vector fields tangent to hyperplane arrangements (in the sequel also referred to as derivation modules) and in particular the question about their freeness over the coordinate ring of the ambient space is a classic topic at the crossroads of discrete geometry and combinatorics, commutative algebra and algebraic geometry initiated in seminal work by K. Saito [S80] and H. Terao [T80a]. Arrangements with free derivation modules are simply called free arrangements. Subsequently, G. Ziegler [Z89] extended the theory by also taking multiplicities into account. Such multiarrangements naturally arise from arrangement constructions, i.e. restrictions to hyperplanes, and the freeness of their mulitderivation modules is closely linked to the freeness of the original simple arrangement due to fundamental work by Ziegler and M. Yoshinaga [Y04].
To decide whether a given (multi-)arrangement is free is a hard problem in general and at its core lies Terao’s conjecture, asserting that freeness only depends on the underlying combinatorial structure provided by the intersection lattice. To this day the conjecture is wide open. Therefore, a natural approach in the investigation of such freeness questions is to restrict the view to special classes of arrangements with additional structure. A pivotal result along this line is a theorem of Terao stating that all (complex) reflection arrangements are free [T80b]. Regarding multiarrangements, another milestone is also due to Terao [T02] establishing the freeness of certain natural multiplicities on Coxeter arrangements. This was subsequently used by Yoshinaga in [Y04] to establish a conjecture of Edelman and Reiner about the freeness of certain deformations of Weyl arrangements.
In the recent work [CK24], M. Cuntz and L. Kühne introduce a new class stemming from simple graphs which are defined as follows.
Definition 1.1 ([CK24, Def. 1.1]).
@@
A natural and important question stemming from the study of free multi-Coxeter arrangements by Terao [T02] is about the existence of constant free multiplicities. Our next result gives the following answer for connected subgraph arrangements.
@@
Let $G$ be a connected graph with at least three vertices and $\mu\equiv k>1$ is a constant multiplicity on ${\mathscr{A}}_{G}$. Then $({\mathscr{A}}_{G},\mu)$ is free if and only if $G$ is a path-graph $P_{n}$ or $G$ is the cycle-graph $C_{3}$ and $k=3$.
To classify free multiplicities in general is a notoriously hard problem. Complete answers are known only in very few cases. E.g. characterizing the free multiplicities even on braid arrangements of rank greater than $3$ is an unresolved problem, cf. [DFMS20]. Nonetheless, in our study it turns out that for a certain subarrangement ${\mathscr{D}}$ of ${\mathscr{A}}_{C_{3}}$ which is obtained from ${\mathscr{A}}_{C_{3}}$ by deleting a single hyperplane, the task is feasible. This constitutes our third theorem.
Theorem 1.4 (Theorem 5.9).
@@
The classification of the free multiplicities on ${\mathscr{D}}$ was previously obtained by M. DiPasquale and M. Wakefield in [DW18] by different methods.
@@
For the notion of inducive freeness, see Definition 2.12.
The paper is organized as follows. In §2 we review the required notions from the theory of free multiarrangements and connected subgraph arrangements. The next section presents the arguments leading to a proof of Theorem 1.2. In §4, we investigate the natural question which connected subgraph arrangements support free constant multiplicities leading to our second main Theorem 1.3. Along the way, in §4.1 we present a non-computational proof of the non-freeness of the connected subgraph arrangements stemming from $G_{1}$ up to $G_{8}$ from Figure 1. This was derived in [CK24] by means of elaborate computer calculations. In our final section we study free multiplicities on two special arrangements, culminating in Theorems 1.4 and 1.5. In the appendices we list some tables containing data which we computed and utilised for our proofs.
@@ 2.1. Hyperplane arrangements, multiarrangements and their freeness
Let $V={\mathbb{K}}^{\ell}$ be an $\ell$-dimensional ${\mathbb{K}}$-vector space. A hyperplane arrangement is a pair $({\mathscr{A}},V)$, where ${\mathscr{A}}$ is a finite collection of hyperplanes in $V$. Usually, we simply write ${\mathscr{A}}$ in place of $({\mathscr{A}},V)$. We write $|{\mathscr{A}}|$ for the number of hyperplanes in ${\mathscr{A}}$. The empty arrangement in $V$ is denoted by $\Phi_{\ell}$.
@@
$D({\mathscr{A}},\mu):=\{\theta\in{\operatorname{Der}}(S)\mid\theta(\alpha_{H})\in\alpha_{H}^{\mu(H)}S\text{ for each }H\in{\mathscr{A}}\}.$
We say that $({\mathscr{A}},\mu)$ is free if $D({\mathscr{A}},\mu)$ is a free $S$-module – a notion introduced by G. Ziegler [Z89, Def. 6], extending K. Saito’s [S80] and H. Terao’s [T80a] freeness for simple arrangements.
@@
and call the pair $({\mathscr{A}}^{H},\kappa)$ the Ziegler-restriction of ${\mathscr{A}}$ to $H$. Ziegler showed the following connection between the freeness of ${\mathscr{A}}$ and the freeness of any of its Ziegler-restrictions.
Theorem 2.2 ([Z89, Thm. 11]).
Suppose ${\mathscr{A}}$ is free with exponents $(1,e_{2},\dots,e_{n})$. Let $H$ be a hyperplane in ${\mathscr{A}}$, then $({\mathscr{A}}^{H},\kappa)$ is free with exponents $(e_{2},\dots,e_{n})$.
@@
Theorem 2.3 ([Z89, Thm. 8], Saito’s criterion).
Let $({\mathscr{A}},\mu)$ be a multiarrangement in $\mathbb{K}^{n}$ and for each $1\leq i\leq n$ let
$\theta_{i}=\sum_{j=1}^{n}p_{ij}\frac{\partial}{\partial x_{j}}\in D({\mathscr{A}},\mu).$
Then $\{\theta_{1},\theta_{2},\dots,\theta_{n}\}$ is a basis for $D({\mathscr{A}},\mu)$ if and only if $\det(p_{ij})=c\cdot Q({\mathscr{A}},\mu)$ for some $c\in\mathbb{K}\setminus\{0\}$.
An arrangement ${\mathscr{A}}$ is called locally free along $H$ for some $H\in{\mathscr{A}}$, if ${\mathscr{A}}_{x}$ is a free arrangement for all $x\in H\backslash\{0\}$. The following result is due to Yoshinaga.
Theorem 2.4 ([Y04, Thm. 2.2]).
Suppose $\ell\geq 3$. Then ${\mathscr{A}}$ is free with exponents $(1,e_{1},\dots,e_{\ell})$ if and only if it is locally free along $H$ for some $H\in{\mathscr{A}}$ and $({\mathscr{A}}^{H},\kappa)$ is free with exponents $(e_{1},\dots,e_{\ell})$.
@@
Suppose $\ell\geq 3$. Let ${\mathscr{A}}$ be locally free along $H\in{\mathscr{A}}$. If ${\mathscr{A}}_{X}$ fails to be free for some $X\in L({\mathscr{A}})$ with $X\not\subset H$, then $({\mathscr{A}}^{H},\kappa)$ is not free.
@@
Recall that an $\ell$-arrangement ${\mathscr{A}}$ with $\operatorname{rank}({\mathscr{A}})=r$ is called generic if $|{\mathscr{A}}_{X}|=k$ for all $X\in L({\mathscr{A}})_{k}$ and $11\leq k\leq r-$ 1. An arrangement ${\mathscr{A}}$ is called totally non-free if $({\mathscr{A}},\mu)$ fails to be free for any multiplicity $\mu\geq\mathbbm{1}$. The following result states that generic arrangements are totally non-free.
Theorem 2.6 ([Y09, Prop. 4.1]).
@@
The following construction from [ATW08] leads to an extension of Terao’s seminal Addition-Deletion Theorem [T80a] for free simple arrangements to multiarrangements.
@@
Let ${\mathscr{A}}^{\prime\prime}=\{H\cap H_{0}\mid H\in{\mathscr{A}}\setminus\{H_{0}\}\ \}$. The Euler multiplicity $\mu^{*}$ of ${\mathscr{A}}^{\prime\prime}$ is defined as follows. Let $Y\in{\mathscr{A}}^{\prime\prime}$. Since the localization ${\mathscr{A}}_{Y}$ is of rank $2$, the multiarrangement $({\mathscr{A}}_{Y},\mu_{Y})$ is free, [Z89, Cor. 7]. According to [ATW08, Prop. 2.1], the module of derivations $D({\mathscr{A}}_{Y},\mu_{Y})$ admits a particular homogeneous basis $\{\theta_{Y},\psi_{Y},D_{3},\ldots,D_{\ell}\}$, such that $\theta_{Y}\notin\alpha_{0}{\operatorname{Der}}(S)$ and $\psi_{Y}\in\alpha_{0}{\operatorname{Der}}(S)$, where $H_{0}=\ker\alpha_{0}$. Then on $Y$ the Euler multiplicity $\mu^{*}$ is defined to be $\mu^{*}(Y)=\operatorname{pdeg}\theta_{Y}$.
@@
It is useful to be able to determine the Euler multiplicity explicitly in some relevant instances. For that purpose, we recall parts of [ATW08, Prop. 4.1].
@@
Theorem 2.10 ([ATW08, Thm. 0.8] Addition-Deletion Theorem for Multiarrangements).
@@
    $({\mathscr{A}},\mu)$ is free with $\exp({\mathscr{A}},\mu)=(b_{1},\ldots,b_{\ell-1},b_{\ell})$;
@@
    $({\mathscr{A}}^{\prime},\mu^{\prime})$ is free with $\exp({\mathscr{A}}^{\prime},\mu^{\prime})=(b_{1},\ldots,b_{\ell-1},b_{\ell}-1)$;
@@
    $({\mathscr{A}}^{\prime\prime},\mu^{*})$ is free with $\exp({\mathscr{A}}^{\prime\prime},\mu^{*})=(b_{1},\ldots,b_{\ell-1})$.
@@
If we restrict to simple multiplicities, Theorem 2.10 recovers Terao’s original Addition-Deletion Theorem from [T80a]. As in the simple case if $({\mathscr{A}},\mu)$ and $({\mathscr{A}}^{\prime},\mu^{\prime})$ are free, then all three statements of Theorem 2.10 hold (see [ATW08, Thm. 0.4]).
@@
As for simple arrangements, if $r({\mathscr{A}})\leq 2$, then $({\mathscr{A}},\mu)$ is inductively free, [Z89, Cor. 7]. Also, ${\mathcal{IFM}}\subseteq{\mathcal{RFM}}$ and ${\mathcal{IFM}}\subseteq{\mathcal{AFM}}$ and all inclusions are strict.
By Remark 2.11, restricting to the subclasses of simple arrangements within the above defined classes, we recover the classes of inductively free ${\mathcal{IF}}$ (cf. [OT92, Def. 4.53]), additively free ${\mathcal{AF}}$ (cf. [A19, Def. 1.6]), and recursively free ${\mathcal{RF}}$ (cf. [OT92, Def. 4.60]) simple arrangements.
@@
Theorem 2.16 ([W07, Thm. 1.5]).
Let ${\mathscr{A}}=\{H_{1},H_{2},H_{3}\}$ be a $2$-arrangement of three lines, $\mu$ a multiplicity on ${\mathscr{A}}$ with $\mu(H_{i})=k_{i},(i=1,2,3)$ and $\exp({\mathscr{A}},\mu)=(d_{1},d_{2})$. Assume that $k_{3}\geq\max\{k_{1},k_{2}\}$ and let $k=k_{1}+k_{2}+k_{3}$.
@@
    $|d_{1}-d_{2}|=\begin{cases}0&\text{if $k$ is even,}\\ 1&\text{if $k$ is odd.}\\ \end{cases}$

### v3

@@ Abstract.
Cuntz and Kühne [5] introduced the class of connected subgraph arrangements ${\mathscr{A}}_{G}$, depending on a graph $G$, and classified all graphs $G$ such that the corresponding arrangement ${\mathscr{A}}_{G}$ is free. We extend their result to the multiarrangement case and classify all graphs $G$ for which the corresponding arrangement ${\mathscr{A}}_{G}$ supports some multiplicity $\mu$ such that the multiarrangement $({\mathscr{A}}_{G},\mu)$ is free.
@@ 1. Introduction
The study of modules of logarithmic vector fields tangent to hyperplane arrangements (in the sequel also referred to as derivation modules) and in particular the question about their freeness over the coordinate ring of the ambient space is a classic topic at the crossroads of discrete geometry and combinatorics, commutative algebra and algebraic geometry initiated in seminal work by K. Saito [12] and H. Terao [13]. Arrangements with free derivation modules are simply called free arrangements. Subsequently, G. Ziegler [19] extended the theory by also taking multiplicities into account. Such multiarrangements naturally arise from arrangement constructions, i.e. restrictions to hyperplanes, and the freeness of their mulitderivation modules is closely linked to the freeness of the original simple arrangement due to fundamental work by Ziegler and M. Yoshinaga [17].
To decide whether a given (multi-)arrangement is free is a hard problem in general and at its core lies Terao’s conjecture, asserting that freeness only depends on the underlying combinatorial structure provided by the intersection lattice. To this day the conjecture is wide open. Therefore, a natural approach in the investigation of such freeness questions is to restrict the view to special classes of arrangements with additional structure. A pivotal result along this line is a theorem of Terao stating that all (complex) reflection arrangements are free [14]. Regarding multiarrangements, another milestone is also due to Terao [15] establishing the freeness of certain natural multiplicities on Coxeter arrangements. This was subsequently used by Yoshinaga in [17] to establish a conjecture of Edelman and Reiner about the freeness of certain deformations of Weyl arrangements.
In the recent work [5], M. Cuntz and L. Kühne introduce a new class stemming from simple graphs which are defined as follows.
Definition 1.1 ([5, Def. 1.1]).
@@
A natural and important question stemming from the study of free multi-Coxeter arrangements by Terao [15] is about the existence of constant free multiplicities. Our next result gives the following answer for connected subgraph arrangements.
@@
Let $G$ be a connected graph with at least three vertices and $\mu\equiv c>1$ is a constant multiplicity on ${\mathscr{A}}_{G}$. Then $({\mathscr{A}}_{G},\mu)$ is free if and only if $G$ is a path-graph $P_{n}$ or $G$ is the cycle-graph $C_{3}$ and $c=3$.
To classify free multiplicities in general is a notoriously hard problem. Complete answers are known only in very few cases. E.g. characterizing the free multiplicities even on braid arrangements of rank greater than $3$ is an unresolved problem, cf. [6]. Nonetheless, in our study it turns out that for a certain subarrangement $\mathscr{D}$} of ${\mathscr{A}}_{C_{3}}$ which is obtained from ${\mathscr{A}}_{C_{3}}$ by deleting a single hyperplane, the task is feasible. This constitutes our third theorem (Theorem 5.9) which was previously obtained by M. DiPasquale and M. Wakefield in [7] by different methods.
Theorem 1.4 ([7]).
@@
For the notion of inductive freeness, see Definition 2.12.
The paper is organized as follows. In §2 we review the required notions from the theory of free multiarrangements and connected subgraph arrangements. The next section presents the arguments leading to a proof of Theorem 1.2. In §4, we investigate the natural question which connected subgraph arrangements support free constant multiplicities leading to our second main Theorem 1.3. Along the way, in §4.1 we present a non-computational proof of the non-freeness of the connected subgraph arrangements stemming from $G_{1}$ up to $G_{8}$ from Figure 1. This was derived in [5] by means of elaborate computer calculations. In our final section we study free multiplicities on two special arrangements, culminating in Theorems 1.4 and 1.5. In the appendices we list some tables containing data which we computed and utilised for our proofs.
@@ 2.1. Hyperplane arrangements, multiarrangements and their freeness
Let $\mathbb{K}$} be a field of characteristic 0 and let $V={\mathbb{K}}^{\ell}$ be an $\ell$-dimensional $\mathbb{K}$}-vector space. A hyperplane arrangement is a pair $({\mathscr{A}},V)$, where $\mathscr{A}$} is a finite collection of hyperplanes in $V$. Usually, we simply write $\mathscr{A}$} in place of $({\mathscr{A}},V)$. We write $|{\mathscr{A}}|$ for the number of hyperplanes in $\mathscr{A}$}. The empty arrangement in $V$ is denoted by $\Phi_{\ell}$.
@@
Let ${\operatorname{Der}}(S)$ be the $S$-module of algebraic $\mathbb{K}$}-derivations of $S$. Using the $\mathbb{Z}$}-grading on $S$, ${\operatorname{Der}}(S)$ becomes a graded $S$-module. For $i=1,\ldots,\ell$, let $D_{i}:=\partial/\partial x_{i}$ be the usual derivation of $S$. Then $D_{1},\ldots,D_{\ell}$ is an $S$-basis of ${\operatorname{Der}}(S)$. We say that $\theta\in{\operatorname{Der}}(S)$ is homogeneous of polynomial degree p provided $\theta=\sum_{i=1}^{\ell}f_{i}D_{i}$, where $f_{i}$ is either 0 or homogeneous of degree $p$ for each $1\leq i\leq\ell$. In this case we write $\operatorname{pdeg}\theta=p$.
@@
[…]
We say that $({\mathscr{A}},\mu)$ is free if $D({\mathscr{A}},\mu)$ is a free $S$-module – a notion introduced by G. Ziegler [19, Def. 6], extending K. Saito’s [12] and H. Terao’s [13] freeness for simple arrangements.
@@
If $\mu=\mathbbm{1}$, we also write $D({\mathscr{A}})$ for $D({\mathscr{A}},\mathbbm{1})$ and say that $\mathscr{A}$} is free if $({\mathscr{A}},\mathbbm{1})$ is. In any case the Euler derivation, $\theta_{E}=\sum_{i=1}^{\ell}x_{i}D_{i}$ belongs to $D({\mathscr{A}})$ and so if $\mathscr{A}$} is free, $1$ is always an exponent of $\mathscr{A}$}, [11, Prop. 4.27].
@@
and call the pair $({\mathscr{A}}^{H},\kappa)$ the Ziegler restriction of $\mathscr{A}$} to $H$. Ziegler showed the following connection between the freeness of $\mathscr{A}$} a simple arrangement and the freeness of any of its Ziegler restrictions.
Theorem 2.2 ([19, Thm. 11]).
Suppose $\mathscr{A}$} is free with exponents $(1,e_{2},\dots,e_{\ell})$. Let $H$ be a hyperplane in $\mathscr{A}$}, then $({\mathscr{A}}^{H},\kappa)$ is free with exponents $(e_{2},\dots,e_{\ell})$.
@@
Theorem 2.3 ([19, Thm. 8], Saito’s criterion).
Let $({\mathscr{A}},\mu)$ be a multiarrangement in $\mathbb{K}^{n}$ and for each $1\leq i\leq\ell$ let
$\theta_{i}=\sum_{j=1}^{\ell}p_{ij}D_{j}\in D({\mathscr{A}},\mu).$
Then $\{\theta_{1},\theta_{2},\dots,\theta_{\ell}\}$ is a basis for $D({\mathscr{A}},\mu)$ if and only if $\det(p_{ij})=c\cdot Q({\mathscr{A}},\mu)$ for some $c\in\mathbb{K}\setminus\{0\}$.
An arrangement $\mathscr{A}$} is called locally free along $H$ for some $H\in{\mathscr{A}}$, if ${\mathscr{A}}_{X}$ is a free arrangement for all subspaces $X$ of $V$ with $\{0\}\neq X\subset H$. The following result is due to Yoshinaga.
Theorem 2.4 ([17, Thm. 2.2]).
Suppose $\mathscr{A}$} is an arrangement in $\mathbb{K}$$^{\ell}$ with $\ell>3$. Then $\mathscr{A}$} is free with exponents $(1=e_{1},e_{2},\dots,e_{\ell})$ if and only if it is locally free along $H$ for some $H\in{\mathscr{A}}$ and $({\mathscr{A}}^{H},\kappa)$ is free with exponents $(e_{2},\dots,e_{\ell})$.
@@
Suppose $\mathscr{A}$} is an arrangement in $\mathbb{K}$$^{\ell}$ with $\ell>3$. Let $\mathscr{A}$} be locally free along $H\in{\mathscr{A}}$. If ${\mathscr{A}}_{X}$ fails to be free for some $X\in L({\mathscr{A}})$ with $X\not\subset H$, then $({\mathscr{A}}^{H},\kappa)$ is not free.
@@
Recall that an $\ell$-arrangement $\mathscr{A}$} with $\operatorname{rank}({\mathscr{A}})=r$ is called generic if $|{\mathscr{A}}_{X}|=k$ for all $X\in L({\mathscr{A}})_{k}$ and $1\leq k\leq r-1$. An arrangement $\mathscr{A}$} is called totally non-free if $({\mathscr{A}},\mu)$ fails to be free for any multiplicity $\mu\geq\mathbbm{1}$. The following result states that generic arrangements are totally non-free.
Theorem 2.6 ([18, Prop. 4.1]).
@@
The following construction from [4] leads to an extension of Terao’s seminal Addition-Deletion Theorem [13] for free simple arrangements to multiarrangements.
@@
Let ${\mathscr{A}}^{\prime\prime}=\{H\cap H_{0}\mid H\in{\mathscr{A}}\setminus\{H_{0}\}\ \}$. The Euler multiplicity $\mu^{*}$ of $\mathscr{A}$$^{\prime\prime}$ is defined as follows. Let $Y\in{\mathscr{A}}^{\prime\prime}$. Since the localization ${\mathscr{A}}_{Y}$ is of rank $2$, the multiarrangement $({\mathscr{A}}_{Y},\mu_{Y})$ is free, [19, Cor. 7]. According to [4, Prop. 2.1], the module of derivations $D({\mathscr{A}}_{Y},\mu_{Y})$ admits a particular homogeneous basis $\{\theta_{Y},\psi_{Y},D_{3},\ldots,D_{\ell}\}$, such that $\theta_{Y}\notin\alpha_{0}{\operatorname{Der}}(S)$ and $\psi_{Y}\in\alpha_{0}{\operatorname{Der}}(S)$, where $H_{0}=\ker\alpha_{0}$. Then on $Y$ the Euler multiplicity $\mu^{*}$ is defined to be $\mu^{*}(Y)=\operatorname{pdeg}\theta_{Y}$.
@@
It is useful to be able to determine the Euler multiplicity explicitly in some relevant instances. For that purpose, we recall parts of [4, Prop. 4.1].
@@
Theorem 2.10 ([4, Thm. 0.8] Addition-Deletion Theorem for Multiarrangements).
@@
    $({\mathscr{A}},\mu)$ is free with $\exp({\mathscr{A}},\mu)=(e_{1},\ldots,e_{\ell-1},b_{\ell})$;
@@
    $({\mathscr{A}}^{\prime},\mu^{\prime})$ is free with $\exp({\mathscr{A}}^{\prime},\mu^{\prime})=(e_{1},\ldots,e_{\ell-1},b_{\ell}-1)$;
@@
    $({\mathscr{A}}^{\prime\prime},\mu^{*})$ is free with $\exp({\mathscr{A}}^{\prime\prime},\mu^{*})=(e_{1},\ldots,e_{\ell-1})$.
@@
If we restrict to simple multiplicities, Theorem 2.10 recovers Terao’s original Addition-Deletion Theorem from [13]. As in the simple case if $({\mathscr{A}},\mu)$ and $({\mathscr{A}}^{\prime},\mu^{\prime})$ are free, then all three statements of Theorem 2.10 hold (see [4, Thm. 0.4]).
@@
As for simple arrangements, if $r({\mathscr{A}})\leq 2$, then $({\mathscr{A}},\mu)$ is inductively free, [19, Cor. 7]. Also, $\mathcal{IFM}$$\subseteq{\mathcal{RFM}}$ and ${\mathcal{IFM}}\subseteq{\mathcal{AFM}}$ and all inclusions are strict.
By Remark 2.11, restricting to the subclasses of simple arrangements within the above defined classes, we recover the classes of inductively free $\mathcal{IF}$} (cf. [11, Def. 4.53]), additively free $\mathcal{AF}$} (cf. [1, Def. 1.6]), and recursively free $\mathcal{RF}$} (cf. [11, Def. 4.60]) simple arrangements.
@@
Theorem 2.16 ([16, Thm. 1.5]).
Let ${\mathscr{A}}=\{H_{1},H_{2},H_{3}\}$ be a $2$-arrangement of three lines, $\mu$ a multiplicity on $\mathscr{A}$} with $\mu(H_{i})=k_{i},(i=1,2,3)$ and $\exp({\mathscr{A}},\mu)=(e_{1},e_{2})$. Assume that $k_{3}\geq\max\{k_{1},k_{2}\}$ and let $k=k_{1}+k_{2}+k_{3}$.
@@
[…]